# Authentic Discrete Diffusion Model

### v1

@@ Abstract
We propose an Authentic Discrete Diffusion (ADD) framework that fundamentally redefines prior pseudo-discrete approaches by preserving core diffusion characteristics directly in the one-hot space through a suite of coordinated mechanisms. Unlike conventional “pseudo” discrete diffusion (PDD) methods, ADD reformulates the diffusion input by directly using float-encoded one-hot class data, without relying on diffusing in the continuous latent spaces or masking policies. At its core, a timestep-conditioned cross-entropy loss is introduced between the diffusion model’s outputs and the original one-hot labels. This synergistic design establishes a bridge between discriminative and generative learning. Our experiments demonstrate that ADD not only achieves superior performance on classification tasks compared to the baseline, but also exhibits excellent text generation capabilities on Image captioning. Extensive ablations validate the measurable gains of each component.
@@ 1 Introduction
Diffusion models have emerged as a powerful class of generative methods, achieving state-of-the-art performance in continuous domains by modeling data generation as a reverse denoising process (Ho et al., 2020). These models operate in continuous-valued vector spaces and are typically trained using mean squared error (MSE) to reconstruct data from Gaussian noise. However, this formulation is inherently incompatible with discrete signals, whose categorical structure and non-Euclidean geometry introduce significant challenges for both optimization stability and sampling fidelity.
Efforts to extend diffusion models to discrete domains have generally followed two unsatisfactory paths. One line of work maps discrete variables into continuous latent spaces via embedding, but these approaches often suffer from degraded generation quality and unstable training (Gong et al., 2023). Another line frames “discrete diffusion” as a masked modeling task, where noise is simulated through random token masking and the model is trained to reconstruct the original input (Google DeepMind, 2025; Nie et al., 2025; Wu et al., 2025; Ye et al., 2025). However, such approaches effectively replicate masked language modeling as popularized by BERT (Devlin et al., 2019), and do not satisfy the formal definition of a diffusion process (Sohl-Dickstein et al., 2015). Moreover, their empirical performance often remains inferior to strong autoregressive baselines. For this reason, we refer to such methods as “pseudo” discrete diffusion (PDD).
In this work, we introduce the Authentic Discrete Diffusion (ADD) model—a framework that preserves the defining properties of diffusion while operating directly in discrete spaces. ADD starts from Gaussian-corrupted one-hot vectors and iteratively denoises them over multiple steps following a well-defined noise schedule. At each step during inference, the model predicts a clean one-hot vector by first applying an $\arg\max$ to its output probabilities, then converting the result into a one-hot representation. This vector is fed back into the next iteration after adding noise with a reduced coefficient. This autoregressive-style feedback loop enables progressive refinement, allowing the model to eliminate uncertainty over time and converge toward semantically precise, categorical outputs.
@@
Figure 1: Comparison between (a) traditional continuous-space diffusion, (b) masking-based pseudo-discrete diffusion (i.e., “pseudo” discrete diffusion (PDD)), and (c) our proposed Authentic Discrete Diffusion (ADD) framework. In (c), the left panel illustrates single-token discrete generation (e.g., classification), while the right panel illustrates multi-token discrete generation (e.g., text generation). Continuous diffusion models operate in Gaussian space, performing noise prediction and MSE-based reconstruction. Pseudo-discrete models mimic diffusion through masked token recovery. In contrast, ADD begins from Gaussian-corrupted one-hot vectors and performs true denoising entirely in the one-hot space. At each step, the model applies an $\arg\max$ to produce discrete predictions, converts them into one-hot vectors, and feeds them into the next iteration after adding noise with a reduced coefficient. This autoregressive-style refinement yields stable and efficient discrete-space diffusion, achieving accurate categorical predictions in only a few steps.
@@
Crucially, ADD applies timestep-conditioned cross-entropy loss to directly enforce correspondence between the predicted and target one-hot vectors, while preventing the diffusion network from over-relying on conditioning at the expense of ignoring the input during training. This avoids the smoothing effects of MSE losses and ensures that outputs respect the mutually exclusive nature of discrete categories. As illustrated in Fig. 1, our method differs fundamentally from prior approaches: continuous diffusion models operate in Gaussian space and perform regression-based noise prediction (Fig. 1a), while pseudo-discrete models simulate diffusion through masked-token recovery (Fig. 1b). In contrast, ADD begins from Gaussian-corrupted one-hot vectors and performs true denoising entirely within the one-hot space, using $\arg\max$-based discretization and iterative noise reduction to progressively refine categorical predictions (Fig. 1c).
We evaluate ADD on two tasks—classification and text generation (e.g., image captioning)—as practical benchmarks for discrete generation. Our contribution lies in establishing a discrete-space diffusion process that is both stable and effective, without relying on diffusion in the embedding space or masking-based approximations. The resulting model achieves strong performance with minimal sampling steps and offers a promising direction for generative modeling in categorical and symbolic domains.
@@
  • We propose ADD, an authentic discrete diffusion framework that preserves the essential principles of generative diffusion while operating entirely in the one-hot space.
  • We introduce a timestep-conditioned cross-entropy loss that directly supervises categorical predictions, avoiding the smoothing effects of MSE and ensuring the mutually exclusive nature of discrete outputs.
  • We design an iterative refinement mechanism that starts from Gaussian-corrupted one-hot vectors and progressively denoises them through $\arg\max$-based discretization and reduced-noise feedback.
  • We evaluate ADD on both classification and text generation (e.g., image captioning) as representative discrete generation tasks, achieving excellent performance with minimal sampling steps, and perform comprehensive ablation studies to isolate the contribution of each component.
@@ 2 Related Work
Diffusion models have achieved state-of-the-art performance in continuous generative modeling, particularly in image synthesis. The foundational Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) models data generation as the reverse of a Gaussian noising process, trained with a mean squared error (MSE) loss to predict added noise. Numerous extensions have enhanced efficiency and fidelity: DDIM (Song et al., 2021) accelerates generation via non-Markovian deterministic sampling; ADM (Dhariwal and Nichol, 2021) incorporates classifier guidance and adversarial techniques; and LDM (Rombach et al., 2022) reduces computation through latent-space diffusion. Despite their success, these methods inherently rely on continuous signal representations and quadratic losses, making them poorly suited for discrete data whose outputs lie on a simplex. One line of work attempts to address this by mapping discrete variables into continuous latent spaces via embedding (Gong et al., 2023). However, such approaches often suffer from degraded generation quality and unstable training, as the continuous diffusion process struggles to model sharp, mutually exclusive distributions.
Pseudo-Discrete Diffusion Models.
Extending diffusion models to discrete domains remains an open challenge. A common alternative reframes “discrete diffusion” as a masked token recovery task (Google DeepMind, 2025; Nie et al., 2025; Wu et al., 2025; Ye et al., 2025). In this view, noise is simulated through random masking of tokens, and the model is trained to reconstruct the original input, effectively mirroring masked language modeling as popularized by BERT (Devlin et al., 2019). While such models can be effective in certain language or multimodal applications, they deviate from the formal definition of diffusion (Sohl-Dickstein et al., 2015), lacking a principled forward noising process and stochastic transition dynamics. Our proposed ADD framework addresses this conceptual gap by preserving the defining characteristics of diffusion in a truly discrete setting. Unlike embedding-based or masking-based approaches, ADD operates directly in one-hot space. In the forward process, a clean one-hot vector is perturbed with Gaussian noise under a variance schedule, producing Gaussian-corrupted one-hot vectors. In the reverse process, the model iteratively denoises these vectors: given a corrupted input, it predicts class probabilities, which are discretized via $\arg\max$ into one-hot form. This estimate is supervised by timestep-conditioned cross-entropy alignment with the target one-hot vector, then re-noised with a reduced coefficient for the next iteration. This $\arg\max$-and-re-noise loop constitutes a genuine diffusion mechanism over discrete symbols, supporting both single-token tasks (e.g., classification) and multi-token generation (e.g., text).
@@
There is growing interest in applying diffusion models or generative models to discriminative tasks such as classification (Wang and Torr, 2022; Wang et al., 2022), segmentation, and object detection. Existing methods often use diffusion indirectly—for example, as a feature extractor (Zhu et al., 2025), a synthetic data generator (Li et al., 2023b; Nguyen et al., 2023; Ma et al., 2023; Wu et al., 2023), or a source of attention cues for zero-shot reasoning (Yang et al., 2025; Liu et al., 2024; Ni et al., 2023). Another line explores zero-shot generative classifiers, where classification is formulated as finding the class condition that minimizes diffusion loss (Li et al., 2023a). While promising, these approaches treat discriminative tasks as secondary outcomes of generative modeling, often requiring external guidance or downstream fine-tuning. Recent works (Chen et al., 2025) also apply diffusion models to segmentation tasks and achieve remarkable performance, but these approaches still operate in continuous space and rely on MSE loss. In contrast, ADD directly formulates discrete prediction as diffusion in one-hot space. For single-token tasks such as classification and multi-token tasks such as text generation, denoising itself serves as categorical prediction. This design eliminates the need for handcrafted supervision signals or embedding-space generation, realizing discrete generation through authentic diffusion in symbolic space.
Figure 2: Overview of the proposed framework. (a) Single-token generation for classification. Ground-truth labels are converted into one-hot vectors and perturbed with Gaussian noise during training. The diffusion model iteratively denoises these corrupted vectors back to categorical one-hot outputs through the “to one” operation—implemented as softmax with timestep-conditioned cross-entropy supervision in training and as $\arg\max$ with one-hot vectorization during sampling. Image features provide conditioning signals that guide the denoising process. (b) Conditioning module. The feature extractor encodes the input image into tokens, where some learnable class tokens interacts with image tokens through stacked Transformer layers to yield the conditioning representation $c$, which is injected into diffusion blocks. (c) Multi-token generation for text. ADD extends naturally from single-label classification to sequence generation by applying Gaussian corruption and denoising to each token in a sequence of one-hot vectors. The diffusion blocks predict categorical distributions for all tokens in parallel, enabling efficient iterative refinement of entire sequences into coherent text.
@@ 3 Methodology
The goal of this work is to develop an Authentic Discrete Diffusion (ADD) framework that retains the essential characteristics of diffusion models while operating directly in one-hot space. Unlike prior pseudo-discrete approaches that rely on masked-token recovery or diffusion in embedding space, ADD defines both forward and reverse processes directly on categorical one-hot vectors. At every step, predictions are projected back to the one-hot domain, ensuring categorical consistency and enabling stable discrete-space diffusion.
@@ 3.1 Background: Diffusion Models
Diffusion models corrupt data progressively with Gaussian noise in a forward process and learn to recover the clean signal through a reverse process. Formally, given data $\mathbf{x}_{0}$, the forward distribution is
$q(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}\!\left(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right),$
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ denotes the variance schedule. The reverse process is parameterized as
$p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}\!\big(\mathbf{x}_{t-1};\,\mu_{\theta}(\mathbf{x}_{t},t),\sigma_{t}^{2}\mathbf{I}\big).$
@@
3.2 Authentic Discrete Diffusion
ADD reformulates diffusion over categorical one-hot vectors. Let $\mathbf{y}_{0}\in\{0,1\}^{K}$ denote a one-hot label over $K$ classes, with $\sum_{k=1}^{K}y_{0}^{(k)}=1$. We also experimented with $\lambda$-hot labels, where $\lambda\in(0,1)$, to make the representations more easily perturbed by Gaussian noise. While this variant led to lower training loss, it did not yield noticeable improvements in accuracy. For this reason, we omit detailed results here and leave further exploration of $\lambda$-hot labels to future work.
@@ 3.2 Authentic Discrete Diffusion
$q(\mathbf{y}_{t}\mid\mathbf{y}_{0})=\mathcal{N}\!\big(\mathbf{y}_{t};\,\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0},\,(1-\bar{\alpha}_{t})\mathbf{I}\big),$
@@
$p_{\theta}(\mathbf{y}_{0}\mid\mathbf{y}_{t},c)=\text{Softmax}\!\big(f_{\theta}(\mathbf{y}_{t},t,c)\big),$
@@
$\mathcal{L}_{\text{CE}}=-\mathbb{E}_{t\sim\mathcal{U}[1,T]}\bar{\alpha}_{t}\sum_{k=1}^{K}y_{0}^{(k)}\log p_{\theta}(y_{0}^{(k)}\mid\mathbf{y}_{t},c).$
@@
$\hat{\mathbf{y}}_{0}=\text{onehot}\!\Big(\arg\max_{k}p_{\theta}(y_{0}^{(k)}\mid\mathbf{y}_{t},c)\Big).$
@@
$\mathbf{y}_{t-1}\sim\mathcal{N}\!\left(\sqrt{\alpha_{t-1}}\,\hat{\mathbf{y}}_{0},\,(1-\alpha_{t-1})\mathbf{I}\right).$

### v2

# In-Situ Tweedie Discrete Diffusion Models

@@ Abstract
While diffusion models excel at generating continuous data such as images, adapting them to discrete tasks has relied on indirect approaches that either operate in continuous embedding spaces or use token masking mechanisms, both of which deviate from modeling the true discrete data distribution that can be theoretically guaranteed by Tweedie’s formula. We propose in-situ Tweedie Discrete Diffusion (TDD), a framework that performs Tweedie’s-formula–guaranteed diffusion directly within the discrete one-hot space—hence “in-situ.” Unlike prior methods that diffuse continuous embeddings or mask tokens, TDD directly corrupts one-hot vectors with Gaussian noise and performs iterative denoising through a timestep-conditioned cross-entropy objective rather than mean-squared-error reconstruction. At each denoising step, the model predicts class probabilities, applies argmax to obtain discrete predictions, converts them to one-hot vectors, and feeds them into the next iteration with progressively reduced noise. This process naturally unifies discriminative classification and generative modeling under a single framework. Experiments demonstrate that TDD achieves strong performance on both image classification and text generation tasks, with extensive ablation studies confirming the effectiveness of each design component. Our work establishes a principled approach to discrete diffusion that preserves the core characteristics of diffusion models while operating natively in discrete space.
@@ 1 Introduction
Figure 1: Comparison between (a) traditional continuous-space diffusion, (b) Mask-based Discrete Diffusion (MDD), and (c) our proposed in-situ Tweedie Discrete Diffusion (TDD) framework. In (c), the left panel illustrates single-token discrete generation (e.g., classification), while the right panel illustrates multi-token discrete generation (e.g., text generation). Continuous diffusion models operate in Gaussian space, performing noise prediction and MSE-based reconstruction. Mask-based discrete models mimic diffusion through masked token recovery. In contrast, TDD begins from Gaussian-corrupted one-hot vectors and performs denoising directly in the one-hot space. At each step, the model applies an $\operatorname*{arg\,max}$ to produce discrete predictions, converts them into one-hot vectors, and feeds them into the next iteration after adding noise with a reduced coefficient. This refinement process yields stable and efficient discrete-space diffusion, achieving accurate categorical predictions in only a few steps.
Diffusion models have emerged as a powerful class of generative methods, achieving state-of-the-art performance in continuous domains by modeling data generation as a reverse denoising process (ho2020denoising). These models operate in continuous-valued vector spaces and are typically trained using mean squared error (MSE) to reconstruct data from Gaussian noise. However, this formulation is inherently incompatible with discrete signals, whose categorical structure and non-Euclidean geometry introduce significant challenges for both optimization stability and sampling fidelity.
Efforts to extend diffusion models to discrete domains have generally followed two unsatisfactory paths. The first line of research formulates discrete diffusion as adding discrete noise, such as discrete Gaussian or discrete uniform perturbations (Austin2021StructuredDD; hoogeboom2021argmax; tran2019discrete). Because these methods often yield poor performance, recent studies have shifted toward token masking strategies, where discrete noise is simulated by randomly replacing input tokens with a special mask token, and the model is trained to reconstruct the original sequence (gemini_diffusion2025; nie2025large; wu2025fast; dream2025). However, this paradigm effectively replicates masked language modeling as in BERT (devlin2019bert) rather than performing true diffusion. All methods in this category—whether based on discrete Gaussian, uniform, or masking noise—fail to satisfy Tweedie’s formula, and therefore lack the theoretical guarantee that the learned process faithfully models the true data distribution. As a result, their empirical performance typically remains inferior to strong autoregressive baselines. The second line of research attempts to circumvent the discreteness issue by mapping tokens into continuous latent spaces via learnable embeddings (gong2022diffuseq). While this approach enables the use of standard Gaussian diffusion techniques, it introduces new challenges: since the embedding space must itself be learned rather than fixed, injecting noise into it can lead to shortcut learning and unstable optimization dynamics. Consequently, generation quality and sample fidelity often degrade, especially when applied to complex discrete datasets.
In this work, we introduce the in-situ Tweedie Discrete Diffusion (TDD) model—a framework that preserves the defining properties of diffusion (i.e., Tweedie’s formula) while operating directly in discrete spaces. TDD begins with Gaussian-corrupted one-hot vectors (i.e., “in-situ”) and iteratively denoises them through multiple steps following a well-defined noise schedule. At each inference step, the model predicts a clean one-hot vector by applying an $\operatorname*{arg\,max}$ operation to its output probabilities, converting the result into a one-hot representation, and re-injecting it into the next iteration with progressively reduced noise. This feedback loop enables progressive refinement, allowing the model to reduce uncertainty over time and converge toward semantically precise, categorical outputs.
@@
Crucially, TDD employs a timestep-conditioned cross-entropy loss that directly enforces correspondence between predicted and target one-hot vectors, while mitigating over-reliance on conditioning signals during training. This design avoids the smoothing effects inherent in MSE losses and ensures that outputs follow the mutually exclusive nature of discrete categories. As illustrated in Fig. 1, our method differs fundamentally from prior approaches: continuous diffusion models operate in Gaussian space and perform regression-based noise prediction (Fig. 1a), while mask-based discrete models simulate diffusion through masked-token recovery (Fig. 1b). In contrast, TDD performs Tweedie’s-formula–guaranteed diffusion entirely within the one-hot space, using $\operatorname*{arg\,max}$-based discretization and iterative noise reduction to progressively refine categorical predictions (Fig. 1c).
We evaluate TDD on two representative tasks, classification and text generation (such as image captioning), which serve as practical benchmarks for discrete generative modeling. The model demonstrates strong performance with only a few sampling steps, providing an efficient solution for generation in categorical domains.
@@
  • We propose TDD, a discrete diffusion framework that preserves Tweedie’s formulation while operating entirely within one-hot space-hence “in-situ”.
  • We introduce a timestep-conditioned cross-entropy loss that directly supervises categorical predictions, avoiding the smoothing effects of MSE and enforcing mutual exclusivity among discrete outputs.
  • We design an iterative refinement mechanism that starts from Gaussian-corrupted one-hot vectors and progressively denoises them via $\operatorname*{arg\,max}$-based discretization and reduced-noise feedback.
  • We evaluate TDD on classification and text generation benchmarks, achieving strong performance with minimal sampling steps and providing extensive ablations to isolate component effects.
@@ 2 Related Work
Diffusion models have achieved state-of-the-art performance in continuous generative modeling, particularly in image synthesis. The foundational Denoising Diffusion Probabilistic Model (DDPM) (ho2020denoising) models data generation as the reverse of a Gaussian noising process, trained with a mean squared error (MSE) loss to predict added noise. Numerous extensions have enhanced efficiency and fidelity: DDIM (song2021denoising) accelerates generation via non-Markovian deterministic sampling; ADM (dhariwal2021diffusion) incorporates classifier guidance and adversarial techniques; and LDM (rombach2022highresolution) reduces computation through latent-space diffusion. Despite their success, these methods inherently rely on continuous signal representations and quadratic losses, making them poorly suited for discrete data whose outputs lie on a simplex. One line of work attempts to address this by mapping discrete variables into continuous latent spaces via embedding (gong2022diffuseq). However, such approaches often suffer from degraded generation quality and unstable training, as the continuous diffusion process struggles to model sharp, mutually exclusive distributions.
Masked-Discrete Diffusion Models.
Extending diffusion models to discrete domains remains an open challenge. A common alternative reframes “discrete diffusion” as a masked token recovery task (gemini_diffusion2025; nie2025large; wu2025fast; dream2025). In this view, noise is simulated through random masking of tokens, and the model is trained to reconstruct the original input, effectively mirroring masked language modeling as popularized by BERT (devlin2019bert). While such models can be effective in certain language or multimodal applications, they deviate from the formal definition of diffusion (sohl2015deep), lacking a principled forward noising process and stochastic transition dynamics. Our proposed TDD framework addresses this conceptual gap by preserving the defining characteristics of diffusion in a truly discrete setting. Unlike embedding-based or masking-based approaches, TDD operates directly in one-hot space. In the forward process, a clean one-hot vector is perturbed with Gaussian noise under a variance schedule, producing Gaussian-corrupted one-hot vectors. In the reverse process, the model iteratively denoises these vectors: given a corrupted input, it predicts class probabilities, which are discretized via $\operatorname*{arg\,max}$ into one-hot form. This estimate is supervised by timestep-conditioned cross-entropy alignment with the target one-hot vector, then re-noised with a reduced coefficient for the next iteration. This $\operatorname*{arg\,max}$-and-re-noise loop constitutes a genuine diffusion mechanism over discrete symbols, supporting both single-token tasks (e.g., classification) and multi-token generation (e.g., text).
@@
There is growing interest in applying diffusion models or generative models to discriminative tasks such as classification (wang2022traditional; wang2022semantic), segmentation, and object detection. Existing methods often use diffusion indirectly—for example, as a feature extractor (zhu2024exploring), a synthetic data generator (li2023open; nguyen2023dataset; ma2023diffusionseg; wu2023diffumask), or a source of attention cues for zero-shot reasoning (yang2024exploring; liu2024vgdiffzero; ni2023ref). Another line explores zero-shot generative classifiers, where classification is formulated as finding the class condition that minimizes diffusion loss (li2023your). While promising, these approaches treat discriminative tasks as secondary outcomes of generative modeling, often requiring external guidance or downstream fine-tuning. Recent works (chen2025gs) also apply diffusion models to segmentation tasks and achieve remarkable performance, but these approaches still operate in continuous space and rely on MSE loss. In contrast, TDD directly formulates discrete prediction as diffusion in one-hot space. For single-token tasks such as classification and multi-token tasks such as text generation, denoising itself serves as categorical prediction. This design eliminates the need for handcrafted supervision signals or embedding-space generation, realizing discrete generation through tweedie diffusion in symbolic space.
Figure 2: Overview of the proposed framework. (a) Single-token generation for classification. Ground-truth labels are converted into one-hot vectors and perturbed with Gaussian noise during training. The diffusion model iteratively denoises these corrupted vectors back to categorical one-hot outputs through the “to one” operation—implemented as softmax with timestep-conditioned cross-entropy supervision in training and as $\operatorname*{arg\,max}$ with one-hot vectorization during sampling. Image features provide conditioning signals that guide the denoising process. (b) Conditioning module. The feature extractor encodes the input image into tokens, where some learnable class tokens interacts with image tokens through stacked Transformer layers to yield the conditioning representation $c$, which is injected into diffusion blocks. (c) Multi-token generation for text. TDD extends naturally from single-label classification to sequence generation by applying Gaussian corruption and denoising to each token in a sequence of one-hot vectors. The diffusion blocks predict categorical distributions for all tokens in parallel, enabling efficient iterative refinement of entire sequences into coherent text.
@@ 3 Methodology
The goal of this work is to develop an Tweedie Discrete Diffusion (TDD) framework that retains the essential characteristics of diffusion models while operating directly in one-hot space. Unlike prior masked-discrete approaches that rely on masked-token recovery or diffusion in embedding space, TDD defines both forward and reverse processes directly on categorical one-hot vectors.
@@ 3.1 Background: Diffusion Models
Diffusion models corrupt data progressively with Gaussian noise in a forward process and learn to recover the clean signal through a reverse process. Formally, given data $\mathbf{x}_{0}$, the forward distribution is:
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right),$
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ denotes the variance schedule. The reverse process is parameterized as:
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}\big(\mathbf{x}_{t-1};\,\mu_{\theta}(\mathbf{x}_{t},t),\sigma_{t}^{2}\mathbf{I}\big).$
@@
3.2 Tweedie Discrete Diffusion
TDD reformulates diffusion over categorical one-hot vectors. Let $\mathbf{y}_{0}\in\{0,1\}^{K}$ denote a one-hot label over $K$ classes, with $\sum_{k=1}^{K}y_{0}^{(k)}=1$. We also experimented with $\lambda$-hot labels, where $\lambda\in(0,1)$, to make the representations more easily perturbed by Gaussian noise. While this variant led to lower training loss, it did not yield noticeable improvements in accuracy. For this reason, we omit detailed results here and leave further exploration of $\lambda$-hot labels to future work.
@@ 3.2 Tweedie Discrete Diffusion
$q(\mathbf{y}_{t}|\mathbf{y}_{0})=\mathcal{N}\big(\mathbf{y}_{t};\,\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0},\,(1-\bar{\alpha}_{t})\mathbf{I}\big),$
@@
$p_{\theta}(\mathbf{y}_{0}|\mathbf{y}_{t},c)=\text{Softmax}\big(f_{\theta}(\mathbf{y}_{t},t,c)\big),$
@@
$\mathcal{L}_{\text{CE}}=-\mathbb{E}_{t\sim\mathcal{U}[1,T]}\bar{\alpha}_{t}\sum_{k=1}^{K}y_{0}^{(k)}\log p_{\theta}(y_{0}^{(k)}|\mathbf{y}_{t},c).$
@@
$\hat{\mathbf{y}}_{0}=\text{onehot}\Big(\operatorname*{arg\,max}_{k}p_{\theta}(y_{0}^{(k)}|\mathbf{y}_{t},c)\Big).$
@@
$\mathbf{y}_{t-1}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t-1}}\,\hat{\mathbf{y}}_{0},\,(1-\bar{\alpha}_{t-1})\mathbf{I}\right).$